OpenStudy (anonymous):

a) Determine whether the Mean Value Theorem applies to " f(x) = ln(2x) " on the interval [1,e] b) If so, determine the points that are guaranteed to exist by the MVT

OpenStudy (anonymous):

\(f(x)=\ln(2x)\) must be continuous on \([1,e]\) and differentiable on \((1,e)\). (You can verify that this is indeed the case.) You then want to find the values of \(c\) such that \[f'(c)=\frac{f(b)-f(a)}{b-a}\] where \(a=1\) and \(b=e\). The derivative of \(\ln(2x)\) is \(f'(x)=\dfrac{1}{x}\), so you have to find \(c\) that satisfies the equation, \[\frac{1}{c}=\frac{\ln(2e)-\ln2}{e-1}\]